OpenStudy (mathmath333):

Find all primes such that, \(\large \color{black}{\begin{align} y^2-4\pmod x\equiv 0\quad \text{and}\hspace{.33em}\\~\\ x^2-1\pmod y\equiv 0 \quad \text{where} \quad \{x,y\}\in\mathbb{P}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

well the problem wants all of prime pairs like \((x,y)\) such that \(x|y^2-4\) and \(y|x^2-1\)

OpenStudy (mathmate):

Here are exhaustive brute force results for 2<x,y<1000, but have no idea where to go from here (5,3),(15,7)

OpenStudy (anonymous):

first suppose that \(x \le y\), since \(y|(x-1)(x+1)\) and \(y>x-1\) it follows that \(y|x+1\) and so \(y=x+1\), there is only one pair of consecutive primes, so we get from here one solution and that is \((x,y)=(2,3)\)

OpenStudy (anonymous):

now going with \(x>y\) Oh, I'm late, I must go somewhere, I'll finish this later ;-))

OpenStudy (mathmath333):

@mathmate \(\large \color{black}{ (5,3) \normalsize\text{ is a valid pair but not } \hspace{.33em}\\~\\ (15,7) \normalsize\text{ as } \large 15 \normalsize\text{ is composite. }\hspace{.33em}\\~\\ }\)

OpenStudy (mathmath333):

mukushla's start is nice but, \(\large \color{black}{\begin{align} 3^2-4\pmod 2\cancel{\equiv} 0\hspace{.33em}\\~\\ \end{align}}\) for the pair of \(\large (2,3)\)

OpenStudy (mathmate):

Oops! Thank you @mathmath333